OpenStudy (anonymous):

A block of mass 3.1 kg hangs vertically from a frictionless pulley of mass 5.9 kg and radius 14 cm. Treat the pulley as a disk. Find: a) the acceleration of the block; b) the tension in the rope; c) the speed of the block after it has fallen 55 cm, assuming it started at rest:

OpenStudy (anonymous):

I_disk=1/2 m r^2

OpenStudy (anonymous):



OpenStudy (anonymous):

is m for both pulley and box?

OpenStudy (anonymous):

block feel the effect of gravity F=5.9 * 9.8

OpenStudy (anonymous):

good question , I don't think so but let me think

OpenStudy (anonymous):

ok so 57.82

OpenStudy (anonymous):

\[\tau= F r\] \[\tau= I a\]

OpenStudy (anonymous):

to reduce confusion \[\tau=I \alpha\]

OpenStudy (anonymous):

\[F r= I \alpha\] you know F , r, I ; Find \[\alpha\] once you find \[\alpha\] divide by r to get a

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

wats I